Greenhab Report – May 21st

End of Season GreenHab Report – 20-May-2019

GreenHab Officer: Shannon Rupert

Environmental control:


Shade cloth (80%) will be kept on throughout the summer

AC unit and heater will be turned off for the summer

Average temperatures: N/A

Hours of supplemental light: none

Daily water usage for crops: n/a

Water in Blue Tank— 0 gallons

Time(s) of watering for crops: n/a

Changes to crops: All plants removed for summer

Narrative: By the end of the week I will have the GreenHab shut down for the summer. I am mixing all the remaining garden soil with perlite and prepping it for planting, and then I will leave it for the summer. I added an additional six inches of potting soil (garden soil with perlite) to the herb garden so the soil level is to the top. I plan to put the remaining soil in the two smaller aquaponics tubs for the summer, and we may use them for planting next year or remove them and plant in pots. At any rate, this will be the best way to keep the soil over summer.

Harvest: There was still a thriving herb garden. Some of the herbs and a few remaining tomato plants were removed and given to a friend in town. All of the remaining seeds from this season were also given to her.

Support/supplies needed: In the fall we will need a new GreenHab Book and new seeds. Any new potting mix should be checked carefully so we don’t reintroduce fungus gnats. Garden Safe Multi-purpose Garden Insect Killer worked better than sticky paper and cost much less so should be used in the future should we have another infestation. Based on observations this season, microgreens are not harvested by the crews so it would be better to invest in larger greens and more substantial plants even though they take longer to grow. The scales will be in the ScienceDome but all other supplies will be in the GreenHab. The GreenHab Book will be in my house.

Operations Report – May 21st

Crew 213 Operations Report 21-May-2019

SOL: 3

Name of person filing report: Dana Levin

Non-nominal systems: Curiosity on switch is loose to the point where when you turn the key the entire assembly turns with it. If this is not remedied the eventual consequence may be tearing the on switch wire connections from torsional forces.

Generator (hours run): overnight 2000-0700

Solar— SOC % (Before generator is run at night): 70% (see above)

Diesel Reading – 78%

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons.

Water (auxillary tank) – 0

Water (static tank) – 400 gallons

Auxillary to Static tank transfer– Yes

Gallons transferred: 100

Water in GreenHab – Greenhab ops have closed for the season, Crew 213 will not be performing greenhab ops

Water (loft) – Static to Loft Pump used – yes

Water Meter: 144100

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Personal rover of director used with her permission by CU- Staff

Hours: 1 hour

Beginning charge: 100%

Ending charge: 68%

Currently charging: yes

Spirit rover used: yes

Hours: 2

Beginning charge: 60%

Ending charge: 15%

Currently charging: yes

Opportunity rover used: yes

Hours: 2

Beginning charge: 100%

Ending charge: 60%

Currently charging: yes

Curiosity rover used:

Hours: 2

Beginning charge: 100%

Ending charge: 72%

Currently charging: yes

Notes on rovers: see above about curiosity off switch

ATV’s Used: no

Reason for use: N/A

Oil Added? no

ATV Fuel Used: no

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs are safely secured near directors residence and will not be used by crew 213 for planned nominal operations

HabCar used and why, where? Yes, to obtain water from town.

CrewCar used and why, where? no (in Grand Junction)

General notes and comments:

– Crew 213 had successful simulation of several medical contingencies on EVA today.

Summary of internet: functional

Summary of suits and radios: Suits are all working well. Crew 212 did a fantastic job of maintaining them. Give our numbers, we are using our own radios.

Summary of Hab operations: nominal

Summary of GreenHab operations: none

Summary of ScienceDome operations: Used for course lectures and briefings. All systems nominal

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: None at this time

Questions, concerns and requests to Mission Support: None needed. Thank you so much for your support of our mission.

Copyright © The Mars Society. All rights reserved. | Main Site